OpenStudy (anonymous):

Please Help Im so confused! Jessie and Bob are financing $425,500 to purchase a house. They obtained a 30/8 balloon mortgage at 6.55%. What will their balloon payment be? $379,659.89 $637,130.99 $357,491.39 $380,040.19 The monthly payment is 2700 but not sure where to go from there

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@thomaster

OpenStudy (phi):

I found a complicated formula. First can you find the future value, using FV= PV (1+r)^n ? (or whatever version you have learned)

OpenStudy (anonymous):

I have to find the monthly formula and then do something with balance formula and that where I'm stuck at.

OpenStudy (phi):

Does this look familiar?

OpenStudy (phi):

how many months before the balloon payment is due ?

OpenStudy (anonymous):

I found the balance formula it is...

OpenStudy (anonymous):

Well 8 year until its due

OpenStudy (phi):

Just making sure. 8 years= 96 months. In your equation, what is n*t (I think that is 96 ?)

OpenStudy (anonymous):

n=12 because the months and t=8 so yes its 96

OpenStudy (phi):

and what is m ?

OpenStudy (anonymous):

I think its the months payed off? Im not sure

OpenStudy (anonymous):

Thank you for your help by the way!

OpenStudy (phi):

so m is 96 ? then n*t is 12*30 = 360 ?

OpenStudy (anonymous):

That would make sense I think I'm not sure

OpenStudy (phi):

and what is P in your equation?

OpenStudy (anonymous):

The principal

OpenStudy (anonymous):

The monthly payment multiplied by the monthly payment were made?

OpenStudy (phi):

ok, I assume i is 0.0655/12 (interest per month) = 0.00545833 and 1+i is 1.00545833

OpenStudy (anonymous):

Yes it is

OpenStudy (phi):

did they give you the monthly payment or did you figure it out ?

OpenStudy (phi):

I keep getting a number around 377,xxx which does not match any of your choices.

OpenStudy (anonymous):

I figured it out

OpenStudy (phi):

what did you get ?

OpenStudy (anonymous):

2731.56

OpenStudy (phi):

for the monthly payment ?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

I used your first formula P* (i * (1+i)^360 )/((1+i)^360 - 1) and got this http://www.wolframalpha.com/input/?i=425500*+%28r+*+%281%2Br%29%5E360+%29%2F%28%281%2Br%29%5E360+-+1%29+where+r%3D+0.0655%2F12

OpenStudy (anonymous):

Oh okay I might have rounded wrong or something! Where do I go from here?

OpenStudy (phi):

I am still confused. The formula I found (with google) gives this answer http://www.wolframalpha.com/input/?i=425500*%281%2Br%29%5E96+-+2703.46*%28+%281%2Br%29%5E96+-+1%29%2Fr+where+r%3D+0.0655%2F12

OpenStudy (anonymous):

Me too I dont understand do you think it might be the one closet to it?

OpenStudy (phi):

your formula gives the answer 377598.83 http://www.wolframalpha.com/input/?i=425500*+%28+%281%2Br%29%5E360+-+%281%2Br%29%5E96+%29%2F%28++%281%2Br%29%5E360+-1%29+where+r%3D+0.0655%2F12 (It rounded to 377599) which matches the first formula (close enough)

OpenStudy (phi):

maybe the last payment (the ballon payment) is the amount due plus the regular payment 377598.83 + 2703.46= 380302.29

OpenStudy (anonymous):

well I had the question before and I thought that was the answer but it was wrong so maybe it is the 379,659

OpenStudy (phi):

Yes, I am (pretty) sure it is one of those 2 choices. But I would like to see how you calculate the exact number.

OpenStudy (anonymous):

Me too im taking an online class and it does not explain it well at all.

OpenStudy (phi):

This might be it. The balloon payment is 377598.83 (using your formula) now add on the interest due on that amount for one month 377598.83*1.00545833= 379,659.89 which matches the first choice.

OpenStudy (anonymous):

That makes sense thank you! When I submitted Ill let you know if it was right or wrong. Thank you so much for your help I REALLY appreciate it!

OpenStudy (phi):

can you figure out the balloon part (how to find 377598.83 )? I typed the formula into wolfram, but they may want you to show your own work.

OpenStudy (anonymous):

You just use the formula and fill it in with the numbers we discussed right? If that then yes I understand

OpenStudy (phi):

ok, good luck.

OpenStudy (anonymous):

Thanks again